# Rotationally invariant dynamical lattice regulators for Euclidean quantum field theories

### v1

@@ 4.3 Symanzik effective theory and universality
(ii) Volume-induced (potential) couplings. The volume factor $V(n)$ multiplies the mass and quartic densities, and its expansion produces vertices with no derivatives on $\phi$:
$S^{(1)}_{\phi\xi}\ \supset\ \sum_{n}a^{d}\;\Big(O(\varsigma\xi)(n)\Big)\Big(\tfrac{1}{2}m_{0}^{2}\phi(n)^{2}+\tfrac{\lambda_{0}}{4!}\phi(n)^{4}\Big),\qquad S^{(2)}_{\phi\xi}\ \supset\ \sum_{n}a^{d}\;\Big(O((\varsigma\xi)^{2})(n)\Big)\Big(\phi^{2}+\phi^{4}\Big).$ (4.29)
On a periodic lattice (or with suitable decay), the linear volume variation may be written as a discrete divergence, so schematically
$V(n)=a^{d}\Big(1+\mathrm{div}^{(0)}\xi(n)+O((\varsigma\xi)^{2})\Big),$ (4.30)
and hence the first-order potential contribution contains
$\sum_{n}a^{d}\,\mathrm{div}^{(0)}\xi(n)\,\Big(\tfrac{1}{2}m_{0}^{2}\phi(n)^{2}+\tfrac{\lambda_{0}}{4!}\phi(n)^{4}\Big).$ (4.31)
A discrete integration by parts then gives, for example,
$\sum_{n}\mathrm{div}^{(0)}\xi(n)\,\phi(n)^{4}\;=\;-\sum_{n}\xi(n)\cdot\nabla^{(0)}\!\big(\phi^{4}\big)(n),$ (4.32)
up to boundary terms (which vanish on the torus). Thus even the non-derivative $\phi^{4}$–vertex arising from the volume factor can be viewed as coupling $\xi$ to a lattice derivative of a local density. In momentum space this implies: every $\xi$-leg enters through at least one lattice derivative (a factor of $\hat{p}$), while the $\phi$-dependence may be quadratic or quartic depending on its origin.
Assuming (SR), connected correlators of local geometric observables built from finitely many $\varsigma\xi$, $g_{\mu\nu}$, $V$, etc. decay exponentially in lattice distance, uniformly in $a$. Equivalently, the geometry sector has a finite correlation length in lattice units, hence a mass scale of order $1/a$ in physical units. As a result, integrating out $\xi$ at fixed $a$ produces a local effective action for the field sector $(U,\Phi)$, with a derivative expansion in powers of (pa)(p\,a) at fixed physical external momenta $p$. This is precisely the input needed for the Symanzik description: all induced operators are local and, by $\mathrm{SE}(d)$ covariance of the regulator, transform covariantly under $\mathrm{SO}(d)$ (and are gauge invariant when gauge fields are present). The geometry sector can and will renormalize the bare-to-renormalized map (shifting masses/couplings and wave-function normalizations), and it can generate higher-dimension $\mathrm{SO}(d)$-scalar operators (including higher-point operators such as $\phi^{6},\phi^{8},\dots$), but under (SR) it cannot produce new relevant/marginal operators beyond those allowed by the target continuum symmetries.
@@
At a conceptual level, the conjecture is a direct application of effective field theory to the enlarged lattice system with fields $(x,U,\Phi)$. First, $\mathrm{SE}(d)$ acts linearly on the embedding field $x$ and leaves both the admissible domain $\mathcal{X}_{\mathrm{adm}}$ and the microscopic action invariant; after quotienting/fixing the global zero modes this symmetry is retained as a covariance of the geometry measure. Consequently, any effective action obtained by integrating out $x$ must be globally $\mathrm{SO}(d)$-covariant (and gauge invariant when gauge fields are present), which yields (i).
Second, the substantive input is (SR): exponential clustering of local geometric observables implies that the geometry sector has a finite correlation length in lattice units, hence integrating out $x$ at fixed cutoff produces a local effective action for $(U,\Phi)$ with a derivative expansion in powers of $pa$ at fixed physical external momenta. This is exactly the locality hypothesis needed for the Symanzik EFT, (4.24). Locality also implies that the geometry sector can only (a) renormalize the bare-to-renormalized map for the operators already present in the target continuum theory and (b) generate higher-dimension $\mathrm{SO}(d)$-covariant operators suppressed by powers of $a$. In particular, absent additional light degrees of freedom in the geometry sector, no new relevant or marginal directions can appear beyond those allowed by the target continuum symmetries, which is (ii). Finally, standard RG universality then gives (iii): once the usual relevant parameters are tuned, different local regulators with the same symmetry content flow to the same continuum fixed point and define the same continuum QFT.
@@ 4.4 Example: one–loop $\beta$–function in scalar $\phi^{4}$ theory
$V(\phi)=\tfrac{1}{2}m_{0}^{2}\phi^{2}+\tfrac{\lambda_{0}}{4!}\phi^{4},$
discretised on the dynamical mesh as in Section 3. On the regular embedding $x_{\rm reg}(n)=an$ this reduces to the standard nearest–neighbour lattice discretisation.
Write
$x(n)=an+\xi(n),$
and assume the short–range geometry hypothesis (SR) from §4.3: connected correlators of local geometric observables built from finitely many $e_{\mu}(n),g_{\mu\nu}(n),V(n)$ decay exponentially in lattice distance, uniformly in $a$. Equivalently (and this is how it is used perturbatively), the $\xi$–sector has a finite correlation length in lattice units. In physical units this means the geometry modes are heavy (their characteristic mass scale is $\sim 1/a$), so integrating them out produces a local effective lattice theory for $\phi$ whose induced interactions are short–range and analytic in the external momenta at scales $\mu\ll 1/a$.
As discussed previously, expanding the scalar action $S_{\phi}[x,\phi]$ in powers of $\xi$ produces two conceptually different classes of geometry–matter couplings.
@@
    $(\nabla\xi)\,(\nabla\phi)(\nabla\phi),\qquad(\nabla\xi)^{2}\,(\nabla\phi)(\nabla\phi),\ \ldots$
    These are the vertices that contribute to wave–function renormalisation and to higher–derivative Symanzik operators.
@@
    $\sum_{n}V(n)\Bigl(\tfrac{1}{2}m_{0}^{2}\phi(n)^{2}+\tfrac{\lambda_{0}}{4!}\phi(n)^{4}\Bigr)$
    produces vertices with no derivatives on the scalar legs, e.g.
    $(\nabla\!\cdot\xi)\,\phi^{2},\qquad(\nabla\!\cdot\xi)\,\phi^{4},\qquad(\nabla\xi)^{2}\,\phi^{2},\qquad(\nabla\xi)^{2}\,\phi^{4},\ \ldots$
    What matters for universality is that under (SR) these vertices describe couplings to a heavy, short–range field.
The one–loop renormalisation of the quartic coupling is determined by the coefficient of the logarithmic dependence on the renormalisation scale $\mu$ in the amputated 1PI four–point function. This coefficient is universal: it arises from the loop–momentum window above $\mu$ but below the cutoff (heuristically $\mu\ll|k|\ll 1/a$), and is insensitive to finite, scheme–dependent redefinitions of bare parameters.
@@
$I_{\rm lat}(a,m_{0}):=\int_{B}\frac{d^{4}k}{(2\pi)^{4}}\,\frac{1}{(\hat{k}^{2}+m_{0}^{2})^{2}}=\frac{1}{16\pi^{2}}\log\frac{1}{a^{2}m_{0}^{2}}+\text{finite},$
and summing the three channels produces the familiar one–loop coefficient $3/(16\pi^{2})$ in the $\phi^{4}$ channel.
@@
(c) Rainbow self–energy from kinetic $\xi\phi\phi$ vertices.
Figure 1: Building blocks in the perturbative universality argument. Solid: $\phi$. Dashed: $\xi$.
(B) Diagrams with internal geometry lines. At one loop, any 1PI four–point contribution that involves at least one internal $\xi$ propagator is controlled by (SR). Indeed, (SR) implies the $\xi$ correlator decays exponentially in lattice units, so the geometry is massive in lattice units. Consequently, at external scales $\mu\ll 1/a$, graphs with internal $\xi$ lines generate local contributions analytic in the external momenta:
  • Kinetic vertices. Kinetic vertices necessarily carry external momenta: already at tree level, two $\xi\phi\phi$ vertices connected by a single internal $\xi$ line produce a momentum–dependent 4–point amplitude (Fig. 1(b)), whose low–energy expansion generates higher–derivative operators in the Symanzik action, not a momentum–independent shift of the $\phi^{4}$ coupling. The corresponding 2–point correction is the rainbow self–energy in Fig. 1(c).
  • Volume vertices. These can and do renormalise the bare–to–renormalised map: integrating out short–range volume fluctuations shifts the coefficients of operators already present in the target theory (mass term, kinetic normalisation, and the quartic coupling) by finite, scheme–dependent amounts (Fig. 2(a)). In addition, exchanging a short–range $\xi$ line between two local volume insertions produces higher multi–field operators; for example Fig. 2(b) corresponds to an induced $\phi^{8}$ interaction, which is irrelevant in $d=4$ and does not affect the running of $\lambda$. Neither mechanism can generate a new logarithmic $\mu$–dependence in the $\phi^{4}$ channel at one loop.
(a) $\xi$-loop dressing (finite matching).
(b) $\xi$-exchange $\Rightarrow\phi^{8}$ (irrelevant).
@@
Assume (SR), so the $\xi$–propagator has a mass scale of order one in lattice units. At external scales $\mu\ll 1/a$, any 1PI graph with at least one internal $\xi$ line produces a local contribution analytic in the external momenta. Consequently, kinetic vertices contribute only to wave–function renormalisation and higher–derivative operators, while volume vertices can shift the local $\phi^{4}$ coefficient by a finite (scheme–dependent) amount and can induce irrelevant multi–field operators (e.g. Fig. 2(b)), but cannot generate an additional $\log\mu$ term in the $\phi^{4}$ channel at one loop. Therefore the coefficient of the logarithmic divergence in the $\phi^{4}$ channel is entirely due to the pure scalar bubbles and equals $3/(16\pi^{2})$, as stated in (4.33); after the usual field renormalisation the renormalised quartic coupling $\lambda(\mu)$ obeys the same one–loop renormalisation–group equation as in the continuum and in the static–lattice regularisation,
@@ 5 Numerical experiments
This section describes the Monte Carlo setup and the numerical diagnostics used to sanity–check the dynamical–lattice regulator (DLR) in a simple two–dimensional scalar test theory. Our aim is a proof–of–concept: to demonstrate that the coupled geometry–matter system can be simulated reliably with standard lattice methods, and that the resulting ensembles exhibit the qualitative properties required of a regulator.
A central working assumption of the dynamical–lattice regulator is that connected correlators of local geometric observables have a finite correlation length in lattice units, uniformly in $L$ (“short–range geometry” or “SR” hypothesis). In the numerics below we treat this as an explicit diagnostic requirement: whenever the matter sector is tuned close to criticality, the geometry sector should remain massive and exhibit exponentially decaying connected correlations.
@@ 5.1 Simulation protocol
with $x$ restricted to the admissible set $\mathcal{X}_{\mathrm{adm}}$ of §2.2. The geometry action $S_{x}$ is the local, E(2)E(2)–invariant action of §2.3, while the scalar action $S_{\phi}$ is the general scalar action of §3.1 evaluated on the same embedded mesh, with the local potential $W_{\mathrm{loc}}$ chosen to give a $\phi^{4}$ theory.
@@
where $\ell_{\mu}$ are the two edge lengths, $A$ is the cell area, $\vartheta$ is the corner angle between the two edges, and $\mathrm{aspect}=\max(\ell_{1},\ell_{2})/\min(\ell_{1},\ell_{2})$. Moves that violate any bound on any affected cell are rejected immediately. In practice these hard constraints prevent near–degenerate cells and keep the simulation in the same connected component as the regular embedding.
@@
  • Scalar move: pick $n\in\Lambda$ and propose $\phi^{\prime}(n)=\phi(n)+\varepsilon_{\phi}$ with $\varepsilon_{\phi}\sim\mathrm{Unif}[-\Delta_{\omega},\Delta_{\omega}]$. Only the four adjacent cells change; we accept with probability $\min\{1,\exp(-\Delta S_{\phi})\}$.
One sweep consists of $L^{2}$ such local proposals. Proposal scales are tuned to give $O(1)$ acceptance; in practice we obtain typical geometry acceptance rates of 70–80% in the parameter range shown.
@@
Run lengths. After thermalisation of $N_{\rm therm}$ sweeps we generate $N_{\rm prod}$ production sweeps, recording measurements every $k$ sweeps. When we vary $L$ and $\mu^{2}$ for the finite–size scaling plots below, we use the same protocol for each parameter point.

### v2

@@ 4.3 Symanzik effective theory and universality
(ii) Volume-induced (potential) couplings. The volume factor $V(n)$ multiplies the mass and quartic densities. At linear order its variation is a scalar made from $\nabla\eta$ (indeed a discrete divergence), so the expansion produces vertices with no derivatives on $\phi$:
$\begin{split}S^{(1)}_{\phi\eta}&\supset\ \sum_{n}a^{d}(\nabla\cdot\eta)(n)\Big(\tfrac{1}{2}m_{0}^{2}\phi(n)^{2}+\tfrac{\lambda_{0}}{4!}\phi(n)^{4}\Big),\\ S^{(2)}_{\phi\eta}&\supset\ \sum_{n}a^{d}|\nabla\eta(n)|^{2}\Big(\phi^{2}+\phi^{4}\Big).\end{split}$ (4.36)
On the periodic torus (or with suitable decay), the linear volume variation is the trace of the lattice Jacobian, so schematically
$V(n)=a^{d}\Big(1+\nabla\!\cdot\eta(n)+O(|\nabla\eta|^{2})\Big).$ (4.37)
Thus the first-order potential contribution includes, in particular, the quartic term
$\sum_{n}a^{d}\,(\nabla\!\cdot\eta)(n)\,\frac{\lambda_{0}}{4!}\,\phi(n)^{4}.$ (4.38)
Applying (4.33) with $F=\phi^{4}$ gives
$\sum_{n}(\nabla\!\cdot\eta)(n)\,\phi(n)^{4}\;=\;-\sum_{n}\eta(n)\cdot\nabla^{*}\big(\phi^{4}\big)(n),$ (4.39)
on the torus. Hence even the non-derivative $\phi^{4}$ term arising from the volume factor may be rewritten as a coupling of $\eta$ to a lattice derivative of a local density. Equivalently, since $x$ enters the microscopic action only through differences $e_{\mu}(n)$, the fluctuation field has the shift symmetry $\eta\mapsto\eta+c$, so every $\eta$ insertion carries at least one lattice derivative (and therefore a factor of $\hat{p}$ in momentum space).
Assuming (SR), connected correlators of local geometric observables built from finitely many $\nabla\eta$, $g_{\mu\nu}$, $V$, etc. decay exponentially in lattice distance, uniformly in $a$. Equivalently, the geometry sector has a finite correlation length in lattice units, hence a mass scale of order $1/a$ in physical units. As a result, integrating out $\eta$ at fixed $a$ produces a local effective action for the field sector $(U,\Phi)$, with a derivative expansion in powers of $pa$ at fixed physical external momenta $p$. This is precisely the input needed for the Symanzik description: all induced operators are local and, by $\mathrm{SE}(d)$ covariance of the regulator, transform covariantly under $\mathrm{SO}(d)$, and are gauge invariant when gauge fields are present. The geometry sector can and will renormalize the bare-to-renormalized map (shifting masses/couplings and wave-function normalizations), and it can generate higher-dimension $\mathrm{SO}(d)$-scalar operators (including higher-point operators such as $\phi^{6},\phi^{8},\dots$), but under (SR) it cannot produce new relevant/marginal operators beyond those allowed by the target continuum symmetries.
@@
At a conceptual level, Conjecture 4.9 is an application of EFT to the enlarged system $(x,U,\Phi)$. Exact $\mathrm{SE}(d)$ covariance of the microscopic measure implies that any effective action obtained after integrating out $x$ is $\mathrm{SO}(d)$-covariant (and gauge invariant when applicable), yielding (i). The substantive input is the short-range geometry hypothesis (SR): exponential clustering in the geometry sector gives a finite correlation length in lattice units, so integrating out $x$ at fixed cutoff produces a local effective action for $(U,\Phi)$ admitting a derivative expansion in powers of $pa$. Consequently the geometry sector can only renormalize the matching of operators already present in the target theory and generate higher-dimension $\mathrm{SO}(d)$-covariant operators suppressed by powers of $a$, giving (ii); standard RG universality then yields (iii) after tuning the usual relevant parameters. Conjecture 4.9 is intended as a concrete target for future work; establishing or falsifying it beyond one loop and in $4d$ gauge theories would be particularly valuable.
@@ 4.4 Example: one–loop $\beta$–function in scalar $\phi^{4}$ theory
$V(\phi)=\tfrac{1}{2}m_{0}^{2}\phi^{2}+\tfrac{\lambda_{0}}{4!}\phi^{4},$ (4.40)
discretised on the dynamical mesh as in Section 3. On the regular embedding $x_{\rm reg}(n)=an$ this reduces to the standard nearest–neighbour lattice discretisation. Write
$x(n)=an+\eta(n),$ (4.41)
and assume the short-range geometry hypothesis (SR) from §4.3: connected correlators of local geometric observables built from finitely many $e_{\mu}(n)$, $g_{\mu\nu}(n)$, $V(n)$ decay exponentially in lattice distance, uniformly in $a$. Equivalently (and this is how it is used perturbatively), the $\eta$–sector has a finite correlation length in lattice units. In physical units this means the geometry modes are heavy with characteristic mass scale $\sim 1/a$, so integrating them out produces a local effective lattice theory for $\phi$ whose induced interactions are short-range and analytic in the external momenta at scales $\mu\ll 1/a$.
As discussed previously, expanding the scalar action $S_{\phi}[x,\phi]$ in powers of $\nabla\eta$ produces two conceptually different classes of geometry–matter couplings.
@@
    $(\nabla\eta)\,(\nabla\phi)(\nabla\phi),\qquad|\nabla\eta|^{2}\,(\nabla\phi)(\nabla\phi),\ \ldots$ (4.42)
    These are the vertices that contribute to wave function renormalisation and to higher-derivative Symanzik operators.
@@
    $\sum_{n}V(n)\Bigl(\tfrac{1}{2}m_{0}^{2}\phi(n)^{2}+\tfrac{\lambda_{0}}{4!}\phi(n)^{4}\Bigr)$ (4.43)
    produces vertices with no derivatives on the scalar legs (though every $\eta$ insertion carries at least one lattice derivative), e.g.
    $(\nabla\!\cdot\eta)\,\phi^{2},\qquad(\nabla\!\cdot\eta)\,\phi^{4},\qquad|\nabla\eta|^{2}\,\phi^{2},\qquad|\nabla\eta|^{2}\,\phi^{4},\ \ldots$ (4.44)
    What matters for universality is that under (SR) these vertices describe couplings to a heavy, short-range field.
The one loop renormalisation of the quartic coupling is determined by the coefficient of the logarithmic dependence on the renormalisation scale $\mu$ in the amputated 1PI four–point function. This coefficient is universal: it arises from the loop–momentum window above $\mu$ but below the cutoff (heuristically $\mu\ll|k|\ll 1/a$), and is insensitive to finite, scheme–dependent redefinitions of bare parameters.
@@
$I_{\rm lat}(a,m_{0})=\int_{B}\frac{d^{4}k}{(2\pi)^{4}}\,\frac{1}{(\hat{k}^{2}+m_{0}^{2})^{2}}=\frac{1}{16\pi^{2}}\log\frac{1}{a^{2}m_{0}^{2}}+\text{finite},$ (4.45)
and summing the three channels produces the familiar one–loop coefficient $3/(16\pi^{2})$ in the $\phi^{4}$ channel. Here $B=[-\pi/a,\pi/a]^{4}$ is the Brillouin zone and
$\hat{k}_{\mu}=\frac{2}{a}\sin\!\Big(\frac{ak_{\mu}}{2}\Big),\qquad\hat{k}^{2}=\sum_{\mu=1}^{4}\hat{k}_{\mu}^{2},$ (4.46)
so that the free lattice propagator is $(\hat{k}^{2}+m_{0}^{2})^{-1}$.
@@
(c) Rainbow self–energy from kinetic $\eta\phi\phi$ vertices.
Figure 1: Building blocks in the perturbative universality argument. Solid: $\phi$. Dashed: $\eta$.
(B) Diagrams with internal geometry lines. At one loop, any 1PI four–point contribution that involves at least one internal $\eta$ propagator is controlled by (SR). Indeed, (SR) implies the $\eta$ correlator decays exponentially in lattice units, so the geometry is massive in lattice units. Consequently, at external scales $\mu\ll 1/a$, graphs with internal $\eta$ lines generate local contributions analytic in the external momenta:
  • Kinetic vertices. Kinetic vertices necessarily carry external momenta: already at tree level, two $\eta\phi\phi$ vertices connected by a single internal $\eta$ line produce a momentum–dependent 4–point amplitude (Fig. 1(b)), whose low–energy expansion generates higher–derivative operators in the Symanzik action, not a momentum–independent shift of the $\phi^{4}$ coupling. The corresponding 2–point correction is the rainbow self–energy in Fig. 1(c).
  • Volume vertices. These can and do renormalise the bare–to–renormalised map: integrating out short–range volume fluctuations shifts the coefficients of operators already present in the target theory (mass term, kinetic normalisation, and the quartic coupling) by finite, scheme–dependent amounts (Fig. 2(a)). In addition, exchanging a short–range $\eta$ line between two local volume insertions produces higher multi–field operators; for example Fig. 2(b) corresponds to an induced $\phi^{8}$ interaction, which is irrelevant in $d=4$ and does not affect the running of $\lambda$. Neither mechanism can generate a new logarithmic $\mu$–dependence in the $\phi^{4}$ channel at one loop.
(a) $\eta$-loop dressing (finite matching).
(b) $\eta$-exchange $\Rightarrow\phi^{8}$ (irrelevant).
@@
Assume (SR), so the $\eta$–propagator has a mass scale of order one in lattice units. At external scales $\mu\ll 1/a$, any 1PI graph with at least one internal $\eta$ line produces a local contribution analytic in the external momenta. Consequently, kinetic vertices contribute only to wave–function renormalisation and higher–derivative operators, while volume vertices can shift the local $\phi^{4}$ coefficient by a finite (scheme–dependent) amount and can induce irrelevant multi–field operators (e.g. Fig. 2(b)), but cannot generate an additional $\log\mu$ term in the $\phi^{4}$ channel at one loop. Therefore the coefficient of the logarithmic divergence in the $\phi^{4}$ channel is entirely due to the pure scalar bubbles and equals $3/(16\pi^{2})$, as stated in (4.47); after the usual field renormalisation the renormalised quartic coupling $\lambda(\mu)$ obeys the same one–loop renormalisation–group equation as in the continuum and in the static–lattice regularisation,
@@ 5 Numerical experiments
This section describes the Monte Carlo setup and numerical diagnostics used to test the dynamical–lattice regulator (DLR) in a two–dimensional scalar theory. The goal is proof–of–concept rather than high–precision spectroscopy: we check (i) algorithmic viability under standard local updates, (ii) that the geometry sector remains massive near matter criticality (the short–range geometry hypothesis (SR)), (iii) suppression of direction–dependent cutoff artefacts at finite lattice spacing, and (iv) preservation of long–distance universality in a standard finite–size scaling test.
A central working assumption is the short–range geometry hypothesis (SR). In the numerics below we treat (SR) as a stringent diagnostic: even when the scalar is tuned close to criticality, the geometry should not develop long range modes.
@@ 5.1 Simulation protocol
with $x$ restricted to the admissible set $\mathcal{X}_{\mathrm{adm}}$ of §2.2. The geometry action $S_{x}$ is the local, E(2)E(2)–invariant action of §2.3, while the scalar action $S_{\phi}$ is taken to be the prototype action $S_{\phi}=S_{\phi}^{\circ}$ of §3.1 (i.e. before $\theta$–averaging), evaluated on the same embedded mesh, with the local potential
$W_{\mathrm{loc}}(\phi)=\frac{1}{2}\mu^{2}\phi^{2}+\lambda\phi^{4}.$ (5.1)
Note that the theory is parametrized by the bare couplings $(\mu^{2},\lambda)$ in lattice units. Our simulation convention for the quartic term is $\lambda\,\phi^{4}$ (no $1/4!$ factor). Thus, relative to the continuum-style normalization $W_{\mathrm{loc}}(\phi)=\tfrac{1}{2}m_{0}^{2}\phi^{2}+\tfrac{\lambda_{0}}{4!}\phi^{4}$ used before, we have
$m_{0}^{2}=\mu^{2},\qquad\lambda_{0}=24\,\lambda.$
We set $\lambda=1$ and vary $\mu^{2}$ in the universality tests, while we use $\lambda=0.1$ and $\mu^{2}=1$ for all other experiments.
@@
This convention fixes the fundamental cycles of the physical torus in finite volume. All diagnostics below are phrased in terms of local twisting/mixing and connected correlators, hence are insensitive to the particular representative chosen for the global $\mathrm{SE}(2)$ zero modes.
@@
where $\ell_{\mu}$ are the two edge lengths, $A$ is the cell area, $\vartheta$ is the corner angle between the two edges, and $\mathrm{aspect}=\max(\ell_{1},\ell_{2})/\min(\ell_{1},\ell_{2})$. The specific numerical values are not tuned: they are conservative bounds chosen to exclude near–degenerate cells while allowing broad fluctuations away from the regular embedding. Moves that violate any bound on any affected cell are rejected immediately. In practice these hard constraints prevent near–degenerate cells and keep the simulation in the same connected component as the regular embedding.
@@
  • Scalar move: pick $n\in\Lambda$ and propose $\phi^{\prime}(n)=\phi(n)+\varepsilon_{\phi}$ with $\varepsilon_{\phi}\sim\mathrm{Unif}[-\Delta_{\phi},\Delta_{\phi}]$. Only the four adjacent cells change; we accept with probability $\min\{1,\exp(-\Delta S_{\phi})\}$.
We distinguish a geometry sweep and a scalar sweep, each consisting of $L^{2}$ single–site proposals of the corresponding type. One Monte Carlo cycle comprises one geometry sweep followed by one scalar sweep. Proposal scales are tuned to give $O(1)$ acceptance rates; in the runs shown the geometry acceptance is typically in the few $\times 10\%$ to $\sim 80\%$ range, depending on $(L,\mu^{2})$.
@@
This minimal choice is designed to keep the geometry disordered at the level of local orientations while the hard constraints control degeneracies; it is not intended as an optimized $S_{x}$ for $d=4$.
Run lengths. For each parameter point we thermalise for $N_{\rm therm}$ sweeps (typically $5\times 10^{3}$), then record $N_{\rm prod}$ sweeps (typically $2\times 10^{4}$), storing measurements every $k$ sweeps (typically $k=10$). Each experiment is repeated over a few dozen independent random seeds, and the same protocol is used uniformly across the $(L,\mu^{2})$ scan underlying the finite–size scaling plots.